# MIRAGE: Scaling Test-Time Inference with Parallel Graph-Retrieval-Augmented Reasoning Chains

### v1

@@ Abstract
Large reasoning models (LRMs) have shown significant progress in test-time scaling through chain-of-thought prompting. Current approaches like search-o1 integrate retrieval augmented generation (RAG) into multi-step reasoning processes but rely on a single, linear reasoning chain while incorporating unstructured textual information in a flat, context-agnostic manner. As a result, these approaches can lead to error accumulation throughout the reasoning chain, which significantly limits its effectiveness in medical question-answering (QA) tasks where both accuracy and traceability are critical requirements. To address these challenges, we propose MIRAGE (Multi-chain Inference with Retrieval-Augmented Graph Exploration), a novel test-time scalable reasoning framework that performs dynamic multi-chain inference over structured medical knowledge graphs. Specifically, MIRAGE 1) decomposes complex queries into entity-grounded sub-questions, 2) executes parallel inference chains, 3) retrieves evidence adaptively via neighbor expansion and multi-hop traversal, and 4) integrates answers using cross-chain verification to resolve contradictions. Experiments on three medical QA benchmarks (GenMedGPT-5k, CMCQA, and ExplainCPE) show that MIRAGE consistently outperforms GPT-4o, Tree-of-Thought variants, and other retrieval-augmented baselines in both automatic and human evaluations. Additionally, MIRAGE improves interpretability by generating explicit reasoning chains that trace each factual claim to concrete chains within the knowledge graph, making it well-suited for complex medical reasoning scenarios. The code will be available for further research.
1 Introduction
@@
Furthermore, current test-time scaling methods face a critical challenge in how they expand their knowledge coverage during inference. Existing retrieval-augmented approaches (Li et al. 2025a) typically acquire unstructured textual information and integrate it into the reasoning process in a flat, context-agnostic manner. This approach to knowledge scaling overlooks the inherent structural relationships and semantic hierarchies within domain knowledge, particularly in specialized fields like medicine where understanding often depends on complex inter-entity relationships, causal chains, and hierarchical taxonomies (Fang et al. 2020; Wen, Wang, and Sun 2024). As a result, even when more computational resources are allocated to retrieve additional information, the flat integration of isolated text fragments limits the system’s ability to perform precise multi-hop reasoning and construct coherent knowledge-grounded arguments. This structural knowledge scaling challenge significantly constrains the effectiveness of test-time scaling in knowledge-intensive domains.
To address these fundamental limitations in test-time scaling, we propose Multi-chain Inference with Retrieval-Augmented Graph Exploration (MIRAGE), a novel reasoning framework that reimagines how computational resources are allocated during inference. Rather than scaling through linear chain extension, MIRAGE implements parallel scaling by decomposing complex queries into semantically coherent sub-questions and executing multiple reasoning trajectories simultaneously. This parallel inference paradigm enables more efficient utilization of computational resources while providing natural mechanisms for cross-chain verification and error correction. To address the knowledge scaling challenge, MIRAGE integrates an adaptive retrieval-reasoning loop that interacts with structured domain knowledge graphs, enabling dynamic exploration of semantic relationships through neighbor expansion and multi-hop traversal strategies. By organizing knowledge retrieval around graph structures rather than flat text fragments, MIRAGE can perform more sophisticated reasoning that leverages the inherent hierarchical and relational properties of domain knowledge. Following the evaluation protocol established by MindMap (Wen, Wang, and Sun 2024), we conduct extensive experiments on three medical question answering (QA) benchmarks, demonstrating that MIRAGE consistently outperforms strong baselines including GPT-4o, Tree-of-Thought variants, and other retrieval-augmented approaches in both automatic metrics and human evaluations across all datasets. In summary, the contributions of this paper are as follows:
@@ 2 Related Work
Prompt-based reasoning methods, including Chain-of-Thought (CoT) prompting (Wei et al. 2022) and its variants such as Self-Consistency (Wang et al. 2023), Tree-of-Thought (Yao et al. 2023a), Graph-of-Thought (Besta et al. 2024), and Program-of-Thought (Chen et al. 2023), have substantially improved the reasoning capabilities of large language models (LLMs) without altering model parameters (Chu et al. 2024; Plaat et al. 2024). These methods rely on carefully designed prompts to elicit multi-step, interpretable rationales, and have been extended to diverse tasks and modalities in various domains (Zhang et al. 2024; Wang et al. 2024).
Recent advances in large reasoning models (LRMs) have led to the development of large language models that are explicitly trained to internalize multi-step reasoning processes. Representative models include DeepSeek-R1 (DeepSeek-AI et al. 2025), OpenAI-o1 (OpenAI et al. 2024b), and Qwen-QwQ (Team 2024b). Through reinforcement learning and reasoning-oriented supervision, these models acquire explicit chain-of-thought abilities and demonstrate strong performance on complex benchmarks such as mathematics and coding (Qu et al. 2025). Nevertheless, both prompt-based and model-based reasoning methods are fundamentally limited by knowledge insufficiency and the accumulation of errors along the reasoning chain, which often leads to unreliable or incorrect answers (Chu et al. 2024; Plaat et al. 2024; Qu et al. 2025). Search-o1 (Li et al. 2025a) and its variants (Li et al. 2025b; Guan et al. 2025) integrate agentic search into the o1-like reasoning process of LLMs but follow a monolithic and sequential paradigm that lacks modularity and cross-step coordination. Extensions such as ARise (Zhang et al. 2025) introduce multi-path exploration via tree-based search with dynamic retrieval. However, these methods all fall short in structured path alignment and global consistency verification, limiting their ability to support complex, multi-hop reasoning.
@@
Retrieval-Augmented Generation (RAG) integrates external knowledge retrieval into the generation process to enhance factual accuracy and reduce hallucinations (Lewis et al. 2020). Classic RAG methods typically employ static, single-step retrieval prior to generation, and have been widely adopted in QA, dialogue, and summarization tasks (Izacard and Grave 2021; Guu et al. 2020). To improve retrieval relevance, later studies explored query rewriting (Ma et al. 2023), reranking (Glass et al. 2022), document compression (Shi et al. 2024), and GraphRAG (Edge et al. 2024; Wen, Wang, and Sun 2024). Meanwhile, generation-side efforts included improved conditioning strategies (Izacard et al. 2023) and post-generation provenance verification to ensure factual consistency (Sankararaman et al. 2024).
Despite those RAG methods get great achievement, they remain limited in multi-step reasoning settings, where information needs evolve dynamically and cannot be satisfied by a single retrieval step. Recent agentic RAG frameworks (Li et al. 2025a; Yao et al. 2023b; Nakano et al. 2022; Asai et al. 2024; Zhang et al. 2025) address this by interleaving retrieval and generation in a loop, enabling models to query external sources based on intermediate reasoning states. However, most existing agentic RAG systems continue to operate over unstructured textual corpora, lacking dynamic access to structured or semantically grounded representations of retrieved content. This greatly limits their ability to support fine-grained, multi-hop reasoning, especially in medical domain that require entity-centric, relational, and hierarchical knowledge modeling.
3 Method
3.1 Task Definition
In this work, we focus on leveraging structured knowledge graphs to support multi-chain inference for medical question answering (QA). Specifically, let queryquery be a natural-language medical question posed by a clinician or patient, and let $\mathcal{G}=(\mathcal{E},\mathcal{R})$ denote a medical knowledge graph, where $\mathcal{E}$ is a set of medical entities (e.g., diseases, drugs, symptoms, foods, etc.) and $\mathcal{R}$ is a set of typed relations capturing associations between these entities. We use $\mathcal{G}$ as the primary external knowledge source because its structured format, consisting of deduplicated entities and clearly defined relations, provides concise and consistent factual information. This helps mitigate the ambiguity and redundancy often found in free-form clinical text. The system returns a single natural-language answer $a\in\mathbb{T}^{+}$, where $\mathbb{T}^{+}$ denotes the set of non-empty free-form text strings.
3.2 Overall Framework
@@
For example, given the query “Why do I keep feeling fatigued even after sleeping well?”, the question decomposer first extracts key entities like fatigue and sleep quality, and generates sub-questions grounded in each, such as identifying causes of fatigue or conditions affecting sleep recovery. The evidence retriever then performs multi-hop reasoning over the knowledge graph, retrieving relevant entities like anemia or thyroid disorders. These intermediate results are merged by the answer synthesizer into a coherent explanation. Throughout the process, the coordinator monitors shared state and triggers each component as inputs become ready. Only output the final answer “Based on your symptoms, chronic fatigue syndrome is the most likely cause…”, keeping the response concise and readable.
@@
Following decomposition, the model enters an iterative retrieval-and-reasoning loop, where each stage is guided by structured prompts (see Appendix A.1). The loop consists of three steps: identifying retrieval targets, querying the knowledge graph, and verbalizing the retrieved evidence. When decoding a sub-question, the model may emit a special search block $\vartheta$, delimited by <|KG_QUERY_BEGIN|> and <|KG_QUERY_END|>, which defines the retrieval target. This block includes one or two entity mentions that are softly matched to entities in $\mathcal{E}$ based on embedding similarity. The matched entity set $\mathcal{E}^{*}$ is defined as follows:
$\mathcal{E}^{*}=\left\{\arg\max_{e\in\mathcal{E}}\mathrm{sim}(\hat{e},e)\;\middle|\;\hat{e}\in\hat{\mathcal{E}},\ \mathrm{sim}(\hat{e},e)\geq\tau\right\}$ (1)
@@
$\mathcal{N}(e)=\left\{(e,r,e^{\prime})\;\middle|\;(e,r,e^{\prime})\in\mathcal{G},\ r\in\mathcal{R}\right\},\quad|\mathcal{N}(e)|\leq k$ (3)
@@
$\mathcal{P}_{h}(e_{1},e_{2})=\left\{p\;\middle|\;p:e_{1}\rightarrow e_{2},\ |p|\leq h\right\}$ (4)
@@
Next, the synthesizer performs pairwise comparisons across all answers, identifying mutually exclusive diagnoses or conflicting therapeutic claims. When such conflicts arise, the system retains the answer whose supporting chain set spans a broader neighborhood of corroborating relations or aligns more closely with the original query. This majority-based verification strategy prefers answers supported by more independently retrieved evidence chains, while suppressing less consistent alternatives. This ensures that the final output is logically coherent and clinically sound before natural-language generation.
@@ 4 Experiment
We evaluate MIRAGE on three public medical QA benchmarks with corresponding knowledge graphs: GenMedGPT-5k with EMCKG, and CMCQA and ExplainCPE with CMCKG (Wen, Wang, and Sun 2024). These datasets cover diverse clinical QA settings, including open-ended questions, multi-turn dialogues, and multiple-choice exams, across both English and Chinese. The paired knowledge graphs contain structured medical facts (e.g., diseases, symptoms, treatments) to support multi-hop reasoning. Dataset statistics are provided in Appendix C.
@@
We compare MIRAGE with a range of baselines covering different reasoning and retrieval strategies. GPT-4o (OpenAI et al. 2024a) serves as a strong general model, and GPT-4o+ToT (Yao et al. 2023a) augments it with Tree-of-Thought prompting for multi-step reasoning. QWQ-32B (Team 2025) represents a large reasoning model trained for end-to-end inference without prompting. We also include retrieval-augmented QWQ-32B variants: BM25 Retriever (Robertson and Zaragoza 2009) (sparse matching), Embedding Retriever (Lewis et al. 2020) (dense similarity), and MindMap (Wen, Wang, and Sun 2024), which uses knowledge graph-based multi-hop retrieval. Search-o1 (Li et al. 2025a) performs dynamic document retrieval with agent-driven iterative refinement.
Implementations Details
@@
Table 1 reports the performance of all methods across three medical QA benchmarks. We summarize the following key observations: (1) MIRAGE consistently achieves the best overall performance in both GPT-4o ranking and answer accuracy. For instance, on GenMedGPT-5k, it reaches a rank of 1.8, and on ExplainCPE, it achieves 84.8% accuracy, outperforming both large-scale models and retrieval-augmented baselines. (2) QWQ-32B outperforms GPT-4o+ToT, indicating that reasoning abilities learned during pretraining may generalize better than prompt-based strategies, even with more powerful LLMs. (3) Static retrieval methods like BM25 and dense retrievers show unstable performance, sometimes helpful but often introducing irrelevant or noisy content. In contrast, structured retrieval via knowledge graphs (e.g., MindMap) provides more consistent improvements, reflecting the need of structure-aware reasoning in medical QA. (4) Search-o1 performs competitively on GenMedGPT-5k and CMCQA but drops notably on ExplainCPE, likely due to noisy web content. MIRAGE remains robust by relying on curated, structured medical knowledge, which enhances both accuracy and reliability.
@@
We explore the impact of the sub-question threshold $N_{q}$ and retrieval threshold $N_{r}$. Specifically, we vary $N_{q}$ while keeping $N_{r}=5$ (its default value), and vary $N_{r}$ while keeping $N_{q}=4$ (its default value). This analysis examines how query decomposition and evidence retrieval jointly affect multi-hop reasoning and answer quality. We report average GPT-4o rank across all datasets and accuracy on ExplainCPE. As shown in Figure 3 (a), performance improves with increasing $N_{q}$ up to a point, then declines as over-decomposition introduces longer or noisier reasoning chains. Although $N_{q}$ only sets an upper bound, higher values often lead to overly aggressive splitting. On dialogue-style queries (e.g., CMCQA), performance may even drop from $N_{q}=1$ to $N_{q}=2$, likely due to disrupted contextual flow. In contrast, Figure 3 (b) shows that increasing $N_{r}$ yields diminishing but generally positive returns, without the sharp decline seen with $N_{q}$. This suggests that the retrieval process is demand-driven: the system often performs fewer retrieval steps than the upper limit $N_{r}$, stopping early when sufficient information is obtained, which helps avoid introducing unnecessary noise or redundancy.
@@
To illustrate the contrast between single-chain and multi-chain graph RAG reasoning, we present a case study comparing Search-o1 and MIRAGE (Figure 5). The baseline Search-o1 performs monolithic reasoning over the entire symptom set using unstructured web retrieval, often leading to information overload and vague explanations. In contrast, MIRAGE identifies toxin exposure and neurological symptoms via separate chains and integrates them to suggest lead poisoning, demonstrating its strength in disentangling complex cases and producing coherent, clinically useful conclusions.
5 Conclusion
In this paper, we proposed MIRAGE, a test-time scalable framework for dynamic multi-chain reasoning over structured medical knowledge graphs. By decomposing complex queries, adaptively structured retrieving evidence, and verifying answers across parallel inference chains, MIRAGE overcomes key limitations of existing linear and unstructured methods. Experiments on three medical QA benchmarks show that MIRAGE consistently improves both performance and interpretability, making it well-suited for high-stakes domains such as medicine, where accurate and traceable reasoning is essential.

### v2

@@ Abstract
Large reasoning models (LRMs) have shown significant progress in test-time scaling through chain-of-thought prompting. Current approaches like search-o1 integrate retrieval augmented generation (RAG) into multi-step reasoning processes but rely on a single, linear reasoning chain while incorporating unstructured textual information in a flat, context-agnostic manner. As a result, these approaches can lead to error accumulation throughout the reasoning chain, which significantly limits its effectiveness in medical question-answering (QA) tasks where both accuracy and traceability are critical requirements. To address these challenges, we propose MIRAGE (Multi-chain Inference with Retrieval-Augmented Graph Exploration), a novel test-time scalable reasoning framework that performs dynamic multi-chain inference over structured medical knowledge graphs. Specifically, MIRAGE 1) decomposes complex queries into entity-grounded sub-questions, 2) executes parallel inference chains, 3) retrieves evidence adaptively via neighbor expansion and multi-hop traversal, and 4) integrates answers using cross-chain verification to resolve contradictions. Experiments on three medical QA benchmarks (GenMedGPT-5k, CMCQA, and ExplainCPE) show that MIRAGE consistently outperforms GPT-4o, Tree-of-Thought variants, and other retrieval-augmented baselines in both automatic and human evaluations. Additionally, MIRAGE improves interpretability by generating explicit reasoning chains that trace each factual claim to concrete chains within the knowledge graph, making it well-suited for complex medical reasoning scenarios.
Codehttps://github.com/Despacitobei/MIRAGE/
Introduction
@@
Furthermore, current test-time scaling methods face a critical challenge in how they expand their knowledge coverage during inference. Existing retrieval-augmented approaches (Li et al. 2025a) typically acquire unstructured textual information and integrate it into the reasoning process in a flat, context-agnostic manner. This approach to knowledge scaling overlooks the inherent structural relationships and semantic hierarchies within domain knowledge, particularly in specialized fields like medicine, where understanding often depends on complex inter-entity relationships, causal chains, and hierarchical taxonomies (Fang et al. 2020; Wen et al. 2024; Wei et al. 2021). As a result, even when more computational resources are allocated to retrieve additional information, the flat integration of isolated text fragments limits the system’s ability to perform precise multi-hop reasoning and construct coherent knowledge-grounded arguments. This structural knowledge scaling challenge significantly constrains the effectiveness of test-time scaling in knowledge-intensive domains.
To address these fundamental limitations in test-time scaling, we propose Multi-chain Inference with Retrieval-Augmented Graph Exploration (MIRAGE), a novel reasoning framework that reimagines how computational resources are allocated during inference. Rather than scaling through linear chain extension, MIRAGE implements parallel scaling by decomposing complex queries into semantically coherent sub-questions and executing multiple reasoning trajectories simultaneously. This parallel inference paradigm enables more efficient utilization of computational resources while providing natural mechanisms for cross-chain verification and error correction. To address the knowledge scaling challenge, MIRAGE integrates an adaptive retrieval-reasoning loop that interacts with structured domain knowledge graphs, enabling dynamic exploration of semantic relationships through neighbor expansion and multi-hop traversal strategies. By organizing knowledge retrieval around graph structures rather than flat text fragments, MIRAGE can perform more sophisticated reasoning that leverages the inherent hierarchical and relational properties of domain knowledge. Following the evaluation protocol established by MindMap (Wen et al. 2024), we conduct extensive experiments on three medical question answering (QA) benchmarks, demonstrating that MIRAGE consistently outperforms strong baselines including GPT-4o, Tree-of-Thought variants, and other retrieval-augmented approaches in both automatic metrics and human evaluations across all datasets. In summary, the contributions of this paper are as follows:
@@
Prompt-based reasoning methods, including Chain-of-Thought (CoT) prompting (Wei et al. 2022), Self-Consistency (Wang et al. 2023), Tree-of-Thought (Yao et al. 2023a), Graph-of-Thought (Besta et al. 2024), and Chain-of-Specificity (Wei et al. 2025b), have substantially improved the reasoning capabilities of large language models (LLMs) without altering model parameters (Chu et al. 2024; Plaat et al. 2024). These methods rely on carefully designed prompts to elicit multi-step, interpretable rationales and have been extended to diverse tasks and modalities in various domains (Zhang et al. 2024).
Recent advances in large reasoning models (LRMs) have led to the development of large language models that are explicitly trained to internalize multi-step reasoning processes. Representative models include DeepSeek-R1 (DeepSeek-AI et al. 2025), OpenAI-o1 (OpenAI et al. 2024b), and Qwen-QwQ (Team 2024b). Through reinforcement learning and reasoning-oriented supervision, these models acquire explicit chain-of-thought abilities and demonstrate strong performance on complex benchmarks such as mathematics and coding (Qu et al. 2025). Nevertheless, both prompt- and model-based approaches are limited by knowledge gaps and error accumulation, often yielding unreliable answers (Chu et al. 2024; Plaat et al. 2024; Qu et al. 2025). Search-o1 (Li et al. 2025a) and its variants (Li et al. 2025b; Guan et al. 2025) integrate agentic search into the o1-like reasoning process of LLMs but follow a monolithic, sequential paradigm lacking modularity and cross-step coordination. ARise (Zhang et al. 2025) adds multi-path tree search with dynamic retrieval but lacks structured path alignment and global consistency, limiting complex multi-hop reasoning.
@@
Retrieval-Augmented Generation (RAG) integrates external knowledge retrieval into the generation process to enhance factual accuracy and reduce hallucinations (Lewis et al. 2020). Classic RAG methods typically employ static, single-step retrieval prior to generation and have been widely adopted in QA, dialogue, and summarization tasks (Guu et al. 2020; Zhu et al. 2025a; Wei et al. 2025a). To improve retrieval relevance, later studies explored query rewriting (Ma et al. 2023), reranking (Glass et al. 2022), document compression (Xu et al. 2024), and GraphRAG (Edge et al. 2024; Wen et al. 2024). Meanwhile, generation-side efforts included improved conditioning strategies (Izacard et al. 2023) and post-generation provenance verification and evaluation (Sankararaman et al. 2024; Zhu et al. 2025b).
Although existing RAG methods have demonstrated significant success, they remain limited in multi-step reasoning settings, where information needs to evolve dynamically and cannot be satisfied by a single retrieval step. Recent agentic RAG frameworks (Li et al. 2025a; Yao et al. 2023b; Nakano et al. 2022; Asai et al. 2024; Zhang et al. 2025) address this by interleaving retrieval and generation in a loop, enabling models to query external sources based on intermediate reasoning states. However, most systems still operate on unstructured text, lacking dynamic access to structured or semantically grounded representations(Wei et al. 2023). This limits their ability to support fine-grained, multi-hop reasoning, especially in the medical domain that requires entity-centric, relational, and hierarchical knowledge modeling.
Method
Task Definition
In this work, we focus on leveraging structured knowledge graphs to support multi-chain inference for medical question answering (QA). Specifically, let queryquery be a natural-language medical question posed by a clinician or patient, and let $\mathcal{G}=(\mathcal{E},\mathcal{R})$ denote a medical knowledge graph, where $\mathcal{E}$ is a set of medical entities (e.g., diseases, drugs, symptoms, foods, …) and $\mathcal{R}$ is a set of typed relations capturing associations between these entities. We use $\mathcal{G}$ as the primary external knowledge source because its structured format, consisting of deduplicated entities and clearly defined relations, provides concise and consistent factual information. This helps mitigate the ambiguity and redundancy often found in free-form clinical text. The system returns a single natural-language answer $a\in\mathbb{T}^{+}$, where $\mathbb{T}^{+}$ denotes the set of non-empty free-form text strings.
Overall Framework
@@
For example, given the query “Why do I keep feeling fatigued even after sleeping well?”, the question decomposer first extracts key entities like fatigue and sleep quality and generates sub-questions grounded in each, such as identifying causes of fatigue or conditions affecting sleep recovery. The evidence retriever then performs multi-hop reasoning over the knowledge graph, retrieving relevant entities like anemia or thyroid disorders. These intermediate results are merged by the answer synthesizer into a coherent explanation. Throughout the process, the coordinator monitors shared state and triggers each component as inputs become ready. Only output the final answer “Based on your symptoms, chronic fatigue syndrome is the most likely cause…”, keeping the response concise and readable.
@@
Following decomposition, the model enters an iterative retrieval-and-reasoning loop, in which each stage is guided by structured prompts (see Appendix A.1). The loop consists of three steps: identifying retrieval targets, querying the knowledge graph, and verbalizing the retrieved evidence. When decoding a sub-question, the model may emit a special search block $\vartheta$, delimited by <|KG_QUERY_BEGIN|> and <|KG_QUERY_END|>, which defines the retrieval target. This block includes one or two entity mentions that are softly matched to entities in $\mathcal{E}$ based on embedding similarity. The matched entity set $\mathcal{E}^{*}$ is defined as follows:
$\mathcal{E}^{*}=\{\arg\max_{e\in\mathcal{E}}\mathrm{sim}(\hat{e},e)\mid\hat{e}\in\hat{\mathcal{E}},\ \mathrm{sim}(\hat{e},e)\geq\tau\}$ (1)
@@
$\mathcal{N}(e){=}\{(e,r,e^{\prime})|(e,r,e^{\prime}){\in}\mathcal{G},r{\in}\mathcal{R}\},|\mathcal{N}(e)|{\leq}k$ (3)
@@
$\mathcal{P}_{h}(e_{1},e_{2})=\{p\mid p:e_{1}\rightarrow e_{2},\ |p|\leq h\}$ (4)
@@
Next, the synthesizer performs pairwise comparisons across all answers, identifying mutually exclusive diagnoses or conflicting therapeutic claims. When such conflicts arise, the system retains the answer whose supporting chain set spans a broader neighborhood of corroborating relations or aligns more closely with the original query. This majority-based verification strategy prefers answers supported by multiple independently retrieved evidence chains, while suppressing less consistent alternatives. This ensures that the final output is logically coherent and clinically sound before natural-language generation.
@@
We evaluate MIRAGE on three public medical QA benchmarks with corresponding knowledge graphs: GenMedGPT-5k with EMCKG, and CMCQA and ExplainCPE with CMCKG (Wen et al. 2024). These datasets cover diverse clinical QA settings, including open-ended questions, multi-turn dialogues, and multiple-choice exams, across both English and Chinese. The paired knowledge graphs contain structured medical facts (e.g., diseases, symptoms, treatments) to support multi-hop reasoning. Dataset statistics are provided in Appendix C.
@@
We compare MIRAGE with a range of baselines covering different reasoning and retrieval strategies. GPT-4o (OpenAI et al. 2024a) serves as a strong general model, and GPT-4o+ToT (Yao et al. 2023a) augments it with Tree-of-Thought prompting for multi-step reasoning. QWQ-32B (Team 2025) represents a large reasoning model trained for end-to-end inference without prompting. We also include retrieval-augmented QWQ-32B variants: BM25 Retriever (Robertson and Zaragoza 2009) (sparse matching), Embedding Retriever (Lewis et al. 2020) (dense similarity), and MindMap (Wen et al. 2024), which uses knowledge graph-based multi-hop retrieval. Search-o1 (Li et al. 2025a) performs dynamic document retrieval with agent-driven iterative refinement.
Implementation Details
@@
Table 1 reports the performance of all methods across three medical QA benchmarks. We summarize the following key observations: (1) MIRAGE consistently achieves the best overall performance in both GPT-4o ranking and answer accuracy. For instance, on GenMedGPT-5k, it reaches a rank of 1.8, and on ExplainCPE, it achieves 84.8% accuracy, outperforming both large-scale models and retrieval-augmented baselines. (2) QWQ-32B outperforms GPT-4o+ToT, indicating that reasoning abilities learned during pretraining may generalize better than prompt-based strategies, even with more powerful LLMs. (3) Static retrieval methods like BM25 and dense retrievers show unstable performance, sometimes helpful but often introducing irrelevant or noisy content. In contrast, structured retrieval via knowledge graphs (e.g., MindMap) provides more consistent improvements, reflecting the need for structure-aware reasoning in medical QA. (4) Search-o1 performs competitively on GenMedGPT-5k and CMCQA but drops notably on ExplainCPE, likely due to noisy web content. MIRAGE remains robust by relying on curated, structured medical knowledge, which enhances both accuracy and reliability.
@@
We explore the impact of the sub-question threshold $N_{q}$ and retrieval threshold $N_{r}$. Specifically, we vary $N_{q}$ while fixing $N_{r}$ at its default of 5, and vary $N_{r}$ while fixing $N_{q}$ at its default of 4. This analysis examines how query decomposition and evidence retrieval jointly affect multi-hop reasoning and answer quality. We report the average GPT-4o rank across all datasets and the accuracy on ExplainCPE. As shown in Figure 3 (a), performance improves with increasing $N_{q}$ up to a point, then declines as over-decomposition introduces longer or noisier reasoning chains. Although $N_{q}$ only sets an upper bound, higher values often lead to overly aggressive splitting. On dialogue-style queries (e.g., CMCQA), performance may even drop from $N_{q}=1$ to $N_{q}=2$, likely due to disrupted contextual flow. In contrast, Figure 3 (b) shows that increasing $N_{r}$ yields diminishing but generally positive returns, without the sharp decline seen with $N_{q}$. This suggests that the retrieval process is demand-driven: the system often performs fewer retrieval steps than the upper limit $N_{r}$, stopping early when sufficient information is obtained, which helps avoid introducing unnecessary noise or redundancy.
@@
To illustrate the contrast between single-chain and multi-chain graph RAG reasoning, we present a case study comparing Search-o1 and MIRAGE (Figure 5). The baseline Search-o1 performs monolithic reasoning over the entire symptom set using unstructured web retrieval, often leading to information overload and vague explanations. In contrast, MIRAGE identifies toxin exposure and neurological symptoms via separate chains and integrates them to suggest lead poisoning, effectively disentangling complex cases to produce coherent, clinically useful conclusions.
Conclusion
In this paper, we propose MIRAGE, a test-time scalable framework for dynamic multi-chain reasoning over structured medical knowledge graphs. By decomposing complex queries, adaptively retrieving structured evidence, and verifying answers across parallel inference chains, MIRAGE overcomes key limitations of existing linear and unstructured methods. Experiments on three medical QA benchmarks show that MIRAGE consistently improves both performance and interpretability, making it well-suited for high-stakes domains such as medicine, where accurate and traceable reasoning is essential.
Acknowledgments
This research was supported by the China National Key R&D Program (Grant No. 2024YFC2707800, 2024YFC2707804), the National Science and Technology Major Project (2025ZD0551300, 2025ZD0551301, 2025ZD0551302), the National Natural Science Foundation of China (62176029, 62506050), the China Postdoctoral Science Foundation Funded Project (2024M763867), and the Chongqing Higher Education Teaching Reform Research Project (No. 242009).